# Steer-Bench: A Benchmark for Evaluating the Steerability of Large Language Models

### v1

@@ Appendix B Data statistics
1 | she, was, relationship, with, friend, did, back, for, it, but | |
2 | age, older, women, gap, young, dating, mature, 30s, olds, attractive | |
4 | single, dating, meet, life, alone, relationships, date, yourself, be, want | |
5 | he, ex, we, relationship, with, ended, back, wasn, time, friend | |
7 | movies, watched, show, anime, characters, horror, tv, scenes, episodes, love | |
8 | attractive, ugly, looks, personality, attraction, appearance, unattractive, beauty, attractiveness, women | |
9 | apps, dating, matches, tinder, bumble, meet, profiles, date, swiping, looking | |
10 | mom, parents, sister, mother, father, family, he, mum, relationship, son | |
12 | income, pay, finances, financially, accounts, rich, women, earns, split, expenses | |
13 | cheese, pizza, eat, meat, add, cooked, salad, delicious, toast, taste | |
15 | friends, female, friendship, friend, platonic, friendships, romantic, attracted, men, relationship | |
16 | men, they, women, are, lack, man, things, woman, all, be | |
17 | compliment, compliments, complimented, complimenting, smile, shirt, say, handsome, told, men | |
18 | smoking, addiction, smoked, cigarettes, nicotine, vape, smokes, cannabis, addict, shrooms | |
19 | straight, bi, lesbian, bisexual, sexuality, lesbians, friends, queer, heterosexual, lgbt | |
@@
1 | feminism, feminists, feminist, equality, rights, movement, women, are, issues, them | |
2 | sub, post, mods, ban, subreddits, on, rights, about, comments, feminist | |
3 | war, military, conscription, drafted, men, combat, selective, wars, are, equality | |
4 | islam, muslim, religions, hijab, muslims, bible, is, islamic, wear, women | |
5 | abortion, abortions, fetus, rights, choice, birth, abort, states, roe, reproductive | |
6 | her, she, was, lawyer, and, with, told, didn, ex, police | |
7 | marriage, divorce, married, alimony, marry, marriages, prenup, relationship, wife, divorced | |
8 | bear, choose, question, forest, woman, encounter, men, alone, be, safer | |
9 | sex, sexual, orgasm, virgin, count, body, women, orgasms, high, having | |
11 | chores, sahm, house, husband, wife, household, housework, job, laundry, tradwife | |
13 | bisexual, lesbians, homophobia, lgbt, are, queer, sexuality, homophobic, gays, friends | |
14 | india, indian, rape, cases, indians, dowry, are, law, women, violence | |
17 | rape, sexual, victims, penetrate, rapists, survey, cdc, data, women, report | |
18 | workers, prostitution, trafficking, industry, prostitutes, is, prostitute, exploitation, job, trafficked | |
@@
1 | trump, vote, party, they, left, biden, voting, democrats, white, election | |
2 | rape, violence, abuse, victims, the, are, and, women, victim, assault | |
3 | this, it, what, re, just, is, comment, post, read, people | |
4 | boys, education, jobs, are, women, teachers, male, is, pay, schools | |
6 | sex, consent, sexual, orgasm, partner, not, women, want, if, can | |
7 | body, fat, beauty, and, people, weight, surgery, appearance, eating, standards | |
8 | military, war, draft, conscription, ukraine, women, drafted, it, are, soldiers | |
9 | characters, character, female, movie, show, it, is, strong, series, woman | |
10 | emotions, emotional, anger, emotion, feelings, men, cry, be, are, can | |
13 | masculinity, masculine, feminine, femininity, positive, traits, be, men, gender, toxic | |
17 | people, bad, it, moral, do, of, be, are, think, morality |
@@
1 | border, immigration, immigrants, asylum, illegals, undocumented, immigrant, deport, deportation, wall | |
2 | gun, guns, shootings, firearms, weapons, amendment, firearm, laws, militia, rifles | |
3 | israel, hamas, palestinians, palestinian, gaza, palestine, genocide, israelis, conflict, hostages | |
4 | religious, god, christianity, christians, bible, religions, commandments, catholic, believe, atheist | |
5 | ukraine, russia, nato, putin, war, ukrainians, russians, crimea, invaded, nuclear | |
6 | schools, education, students, teachers, degree, teacher, debt, funding, teaching, colleges | |
7 | she, vp, has, candidate, think, that, biden, president, of, would | |
8 | healthcare, insurance, medicare, aca, companies, medicaid, government, countries, us, premiums | |
9 | biden, trump, candidate, voters, party, democrats, election, joe, 2020, democratic | |
10 | harris, trump, biden, vote, candidate, voters, win, election, democrats, democratic | |
11 | comment, conversation, op, response, point, reply, understand, argument, thread, discussion | |
12 | taxes, spending, deficit, cut, capital, increase, rates, wealthy, taxing, government | |
13 | covid, vaccines, pandemic, vaccinated, flu, vaccination, deaths, lockdowns, polio, measles | |
14 | news, media, cnn, sources, msnbc, npr, propaganda, journalism, journalists, unbiased | |
15 | capitalism, socialism, communism, socialist, capitalist, marx, marxism, marxist, democracy, economy | |
16 | trans, gender, puberty, dysphoria, transgender, medical, transitioning, hormones, minors, therapy | |
17 | kamala, she, harris, trump, biden, candidate, primary, voters, election, democrats | |
18 | protests, insurrection, riots, riot, rioters, peaceful, protestors, protesters, capital, election | |
19 | mods, banned, conservatives, post, askconservatives, subreddits, rules, liberal, askaliberal, moderation | |
@@
1 | ukraine, russia, war, putin, israel, nato, hamas, iran, are, military | |
2 | court, he, case, is, president, supreme, judge, law, trial, immunity | |
3 | biden, trump, debate, in, president, vote, has, it, joe, who | |
4 | border, immigration, bill, immigrants, asylum, they, mexico, in, is, legal | |
5 | kamala, biden, harris, for, they, has, vote, as, debate, if | |
6 | left, party, democrats, are, right, republicans, conservatives, liberal, conservative, liberals | |
7 | abortion, abortions, birth, women, is, roe, ivf, babies, issue, pregnant | |
8 | blue, state, live, california, cities, states, rural, texas, move, liberal | |
9 | harris, biden, to, campaign, vote, that, if, they, has, win | |
11 | gun, guns, firearms, amendment, shootings, ban, assault, laws, mass, carry | |
12 | christian, religion, christians, bible, catholic, christianity, are, islam, commandments, churches | |
13 | violence, insurrection, protests, riots, capitol, january, antifa, was, blm, riot | |
14 | news, media, cnn, msnbc, the, journalism, propaganda, outlets, bias, journalists | |
15 | polls, polling, pollsters, election, betting, trump, is, voters, data, polled | |
19 | vote, voting, party, election, volunteer, candidate, blue, in, will, are | |
@@
1 | to, is, for, it, as, with, have, woman, just, out | |
2 | inflation, economy, prices, that, tax, for, are, trump, tariffs, all | |
4 | kamala, to, trump, it, harris, as, biden, have, vote, will | |
8 | he, convicted, trump, to, case, trial, judge, jury, prison, court | |
9 | men, trans, are, they, it, gender, black, be, woman, as | |
13 | border, bill, it, immigrants, in, immigration, wall, are, republicans, asylum |
@@
1 | reddit, post, what, re, comments, opinion, google, wrong, downvoted, don | |
2 | discs, copies, buy, license, digitally, games, download, store, sales, internet | |
3 | platinum, trophies, achievements, plat, platinums, platinumed, getting, hours, enjoy, games | |
4 | ssd, storage, 2tb, expansion, hdd, 4tb, heatsink, install, seagate, ps5 | |
5 | xbox, gen, console, games, 360, 4k, disc, upgrade, performance, storage | |
6 | ps2, ps1, playstation, nes, owned, sega, memories, n64, atari, snes | |
7 | xbox, exclusives, microsoft, platform, consoles, sony, market, nintendo, platforms, exclusivity | |
8 | gamepass, month, subscription, xbox, games, conversion, buy, gp, price, sales | |
9 | he, wife, kids, dad, together, likes, gift, play, happy, roblox | |
10 | horizontal, overheating, airflow, cooling, vents, ps5, ventilation, exhaust, cabinet, thermal | |
11 | cod, mw3, mw2, warfare, battlefield, warzone, multiplayer, campaigns, modern, bo2 | |
12 | sale, price, friday, full, discount, buy, 60, deals, waiting, games | |
13 | pc, gaming, ps5, consoles, build, pcs, gpu, performance, windows, steam | |
14 | boss, difficulty, fight, level, difficult, enemies, mode, parry, game, hours | |
15 | female, she, characters, protagonist, gender, yasuke, white, japanese, samurai, censored | |
16 | region, psn, vpn, dlc, accounts, steam, sony, uk, requirement, eu | |
17 | router, wifi, ethernet, network, modem, mbps, dns, 5ghz, hotspot, fiber | |
18 | drift, stick, controllers, controller, sticks, dualsense, fix, left, launch, joystick | |
19 | forza, cars, horizon, turismo, motorsport, motorstorm, arcade, motorfest, racer, sim | |
@@
1 | viktor, lore, characters, show, jayce, zaun, ekko, ionia, episodes, hextech | |
2 | he, worlds, year, player, top, bad, lck, been, performance, mid | |
3 | smurfs, smurfing, accounts, mmr, valve, alt, emerald, ban, matchmaking, level | |
4 | dota, friends, life, fun, games, moba, time, addiction, addicted, much | |
5 | post, read, comments, argument, opinion, downvotes, downvote, discussion, thread, comprehension | |
6 | chat, mute, voice, muting, comms, pinging, toxicity, typing, use, mic | |
10 | he, nerfed, nerfs, kit, lane, champ, mid, ult, buff, patch | |
11 | support, ult, mid, lane, she, adc, lux, jungle, champion, mage | |
12 | fps, gpu, pc, issue, steam, dota, ryzen, reconnect, server, ssd | |
15 | baron, draft, they, pick, dragon, game, geng, call, rumble, drakes | |
17 | reports, 12k, system, toxic, griefing, communication, overwatch, commends, comm, scores | |
18 | mouse, camera, keyboard, hotkeys, buttons, press, cursor, hotkey, screen, bind | |
@@
1 | post, comment, conversation, here, argument, don, op, read, have, reply | |
2 | abortion, fetus, abortions, woman, choice, medical, unborn, conception, is, care | |
4 | wives, roles, equal, church, misogyny, paul, she, feminism, authority, misogynistic | |
5 | trans, gender, transgender, dysphoria, identity, intersex, medical, transitioning, cis, transgenderism | |
6 | israel, hamas, palestinians, palestine, gaza, genocide, israelis, idf, zionism, conflict | |
7 | islam, muslims, muslim, islamic, sharia, religions, islamophobia, europe, quran, iran | |
8 | atheist, school, raised, religion, believed, catholic, parents, religious, became, wasn | |
9 | slavery, slaves, slave, bible, israelites, exodus, leviticus, servitude, enslaved, testament | |
10 | evolution, species, apes, darwin, macroevolution, ancestor, creationism, organisms, scientists, abiogenesis | |
11 | science, scientific, scientists, scientist, religion, hypothesis, theories, einstein, physics, newton | |
12 | catholics, protestants, catholicism, protestant, pope, orthodoxy, apostolic, protestantism, denominations, bishops | |
13 | music, songs, metal, bands, rap, listening, satanic, genre, hymns, taylor | |
14 | rapture, end, tribulation, raptured, earth, matthew, generation, time, angels, soon | |
15 | hell, torment, place, gehenna, heaven, death, hades, eternity, sheol, soul | |
16 | schools, education, teach, teachers, curriculum, oklahoma, commandments, religions, district, classrooms | |
18 | dress, clothing, hijab, modesty, woman, dressing, covering, nakedness, muslim, nudity | |
19 | religion, religions, people, control, society, world, power, organized, we, masses |
@@
1 | israel, hamas, palestinians, palestine, palestinian, israelis, zionist, idf, zionists, zionism | |
2 | marry, marriage, relationship, he, married, muslim, family, convert, want, date | |
3 | she, muslim, relationship, family, tell, but, will, marry, convert, want | |
4 | hijab, women, cover, niqab, clothing, modesty, naked, muslim, she, hijabi | |
5 | homosexuality, lgbt, lgbtq, queer, homosexual, homophobic, straight, are, gays, sin | |
6 | hindus, indian, bangladesh, caste, hindutva, pakistanis, bangladeshi, bjp, bengali, culture | |
7 | woman, rights, islam, feminism, feminist, muslim, misogynistic, misogyny, feminists, is | |
9 | sin, forgive, forgiveness, repent, allah, repentance, forgiven, forgives, sinning, repenting | |
10 | islam, left, leaving, muslim, started, about, it, believe, convert, didn | |
11 | post, troll, account, comments, sub, argument, debate, reply, read, arguments | |
12 | universe, god, existence, exist, infinite, evidence, dependent, believe, argument, creation | |
13 | fasting, ramadan, eat, water, days, fasts, health, weight, up, if | |
14 | alcohol, drink, drinking, addiction, smoking, haram, alcoholic, weed, nicotine, harmful | |
15 | arabic, language, translations, quran, read, speak, learning, qur, tafsir, arab | |
16 | hadith, hadiths, bukhari, quran, scholars, narration, fabricated, books, authenticity, qur | |
@@
1 | phone, media, apps, app, screen, scrolling, reddit, instagram, youtube, tiktok | |
2 | porn, addiction, masturbation, is, that, with, sexual, as, have, watching | |
3 | weight, eat, diet, and, healthy, protein, foods, calorie, body, meals | |
4 | gym, workout, exercise, do, it, week, day, feel, working, like | |
5 | adhd, with, medication, diagnosed, can, meds, like, or, get, be | |
6 | thank, for, sharing, it, post, to, and, words, hear, advice | |
7 | read, books, reading, atomic, life, was, habits, this, changed, helped | |
8 | friends, people, social, with, meet, be, talk, group, like, new | |
9 | weed, smoking, smoke, quit, nicotine, quitting, smoked, years, turkey, vape | |
10 | goals, tasks, task, do, list, it, time, day, work, small | |
11 | life, age, young, old, 20s, of, time, are, 40, 30 | |
13 | her, relationship, he, with, will, but, ex, love, for, time | |
14 | quote, is, man, life, he, be, what, will, quotes, from | |
15 | year, was, myself, of, been, 2024, have, years, when, be | |
16 | study, studying, exam, hours, time, if, break, exams, this, minutes | |
17 | post, comment, this, op, advice, it, re, don, like, understand | |
18 | addiction, addictions, drugs, addicted, with, are, can, but, life, will | |
19 | women, dating, date, are, be, relationship, will, that, girl, rejection | |
@@
1 | her, to, that, the, relationship, is, with, in, do, cheating | |
2 | thank, for, much, appreciate, proud, sharing, all, advice, words, keep | |
3 | phone, media, app, apps, use, tiktok, and, facebook, scrolling, screen | |
6 | ugly, looks, attractive, appearance, and, beauty, women, of, yourself, re | |
10 | parents, family, mom, dad, mother, that, he, their, was, be | |
19 | thoughts, meditation, mind, thought, thinking, brain, intrusive, mindfulness, think, be |
@@
1 | xavi, coach, laporta, manager, season, he, club, team, stay, has | |
2 | ref, var, refs, foul, offside, penalty, referee, referees, negreira, madrid | |
3 | we, half, game, score, chances, goal, team, goals, defense, up | |
4 | ballon, award, vini, won, messi, votes, awards, carvajal, deserved, euros | |
5 | mbappe, mbappé, psg, madrid, kylian, ronaldo, player, will, world, and | |
6 | vini, vinicius, neymar, him, player, brazil, jr, has, but, fans | |
7 | mods, post, comments, sub, reddit, twitter, mod, banned, users, discussion | |
9 | kits, logo, shirt, jerseys, nike, authentic, design, crest, replica, badge | |
10 | pedri, gavi, fdj, role, midfield, midfielder, injured, pivot, play, gundo | |
11 | kroos, modric, toni, luka, modrić, retire, midfield, midfielder, midfielders, season | |
12 | contract, sell, clause, salary, million, pay, fee, transfer, loan, value | |
16 | yamal, messi, age, 17, kid, talent, him, young, player, will | |
18 | madrid, fans, barca, sub, real, support, hate, club, comments, barcelona | |
@@
1 | klay, thompson, warriors, was, but, bench, contract, in, with, you | |
2 | kerr, minutes, klay, steph, coaching, this, vets, lineups, curry, when | |
3 | ad, lebron, center, he, ball, defense, play, can, big, offense | |
6 | contract, value, trade, million, salary, pay, cap, player, option, paid | |
7 | refs, foul, calls, fouls, officiating, ball, fouled, replay, lakers, bounds | |
11 | draft, son, nba, 55th, lebron, bron, his, james, be, lakers | |
12 | reddit, post, read, comments, re, opinion, what, said, internet, downvotes | |
16 | coach, coaches, coaching, fired, hire, hc, nba, riley, assistants, hiring | |
18 | fans, kobe, laker, lakers, lebron, hate, sub, media, haters, hater | |

### v2

@@ Appendix B Data statistics
| 1 | she, was, relationship, with, friend, did, back, for, it, but | |
| 2 | age, older, women, gap, young, dating, mature, 30s, olds, attractive | |
| 4 | single, dating, meet, life, alone, relationships, date, yourself, be, want | |
| 5 | he, ex, we, relationship, with, ended, back, wasn, time, friend | |
| 7 | movies, watched, show, anime, characters, horror, tv, scenes, episodes, love | |
| 8 | attractive, ugly, looks, personality, attraction, appearance, unattractive, beauty, attractiveness, women | |
| 9 | apps, dating, matches, tinder, bumble, meet, profiles, date, swiping, looking | |
| 10 | mom, parents, sister, mother, father, family, he, mum, relationship, son | |
| 12 | income, pay, finances, financially, accounts, rich, women, earns, split, expenses | |
| 13 | cheese, pizza, eat, meat, add, cooked, salad, delicious, toast, taste | |
| 15 | friends, female, friendship, friend, platonic, friendships, romantic, attracted, men, relationship | |
| 16 | men, they, women, are, lack, man, things, woman, all, be | |
| 17 | compliment, compliments, complimented, complimenting, smile, shirt, say, handsome, told, men | |
| 18 | smoking, addiction, smoked, cigarettes, nicotine, vape, smokes, cannabis, addict, shrooms | |
| 19 | straight, bi, lesbian, bisexual, sexuality, lesbians, friends, queer, heterosexual, lgbt | |
@@
| 1 | feminism, feminists, feminist, equality, rights, movement, women, are, issues, them | |
| 2 | sub, post, mods, ban, subreddits, on, rights, about, comments, feminist | |
| 3 | war, military, conscription, drafted, men, combat, selective, wars, are, equality | |
| 4 | islam, muslim, religions, hijab, muslims, bible, is, islamic, wear, women | |
| 5 | abortion, abortions, fetus, rights, choice, birth, abort, states, roe, reproductive | |
| 6 | her, she, was, lawyer, and, with, told, didn, ex, police | |
| 7 | marriage, divorce, married, alimony, marry, marriages, prenup, relationship, wife, divorced | |
| 8 | bear, choose, question, forest, woman, encounter, men, alone, be, safer | |
| 9 | sex, sexual, orgasm, virgin, count, body, women, orgasms, high, having | |
| 11 | chores, sahm, house, husband, wife, household, housework, job, laundry, tradwife | |
| 13 | bisexual, lesbians, homophobia, lgbt, are, queer, sexuality, homophobic, gays, friends | |
| 14 | india, indian, rape, cases, indians, dowry, are, law, women, violence | |
| 17 | rape, sexual, victims, penetrate, rapists, survey, cdc, data, women, report | |
| 18 | workers, prostitution, trafficking, industry, prostitutes, is, prostitute, exploitation, job, trafficked | |
@@
| 1 | trump, vote, party, they, left, biden, voting, democrats, white, election | |
| 2 | rape, violence, abuse, victims, the, are, and, women, victim, assault | |
| 3 | this, it, what, re, just, is, comment, post, read, people | |
| 4 | boys, education, jobs, are, women, teachers, male, is, pay, schools | |
| 6 | sex, consent, sexual, orgasm, partner, not, women, want, if, can | |
| 7 | body, fat, beauty, and, people, weight, surgery, appearance, eating, standards | |
| 8 | military, war, draft, conscription, ukraine, women, drafted, it, are, soldiers | |
| 9 | characters, character, female, movie, show, it, is, strong, series, woman | |
| 10 | emotions, emotional, anger, emotion, feelings, men, cry, be, are, can | |
| 13 | masculinity, masculine, feminine, femininity, positive, traits, be, men, gender, toxic | |
| 17 | people, bad, it, moral, do, of, be, are, think, morality |
@@
| 1 | border, immigration, immigrants, asylum, illegals, undocumented, immigrant, deport, deportation, wall | |
| 2 | gun, guns, shootings, firearms, weapons, amendment, firearm, laws, militia, rifles | |
| 3 | israel, hamas, palestinians, palestinian, gaza, palestine, genocide, israelis, conflict, hostages | |
| 4 | religious, god, christianity, christians, bible, religions, commandments, catholic, believe, atheist | |
| 5 | ukraine, russia, nato, putin, war, ukrainians, russians, crimea, invaded, nuclear | |
| 6 | schools, education, students, teachers, degree, teacher, debt, funding, teaching, colleges | |
| 7 | she, vp, has, candidate, think, that, biden, president, of, would | |
| 8 | healthcare, insurance, medicare, aca, companies, medicaid, government, countries, us, premiums | |
| 9 | biden, trump, candidate, voters, party, democrats, election, joe, 2020, democratic | |
| 10 | harris, trump, biden, vote, candidate, voters, win, election, democrats, democratic | |
| 11 | comment, conversation, op, response, point, reply, understand, argument, thread, discussion | |
| 12 | taxes, spending, deficit, cut, capital, increase, rates, wealthy, taxing, government | |
| 13 | covid, vaccines, pandemic, vaccinated, flu, vaccination, deaths, lockdowns, polio, measles | |
| 14 | news, media, cnn, sources, msnbc, npr, propaganda, journalism, journalists, unbiased | |
| 15 | capitalism, socialism, communism, socialist, capitalist, marx, marxism, marxist, democracy, economy | |
| 16 | trans, gender, puberty, dysphoria, transgender, medical, transitioning, hormones, minors, therapy | |
| 17 | kamala, she, harris, trump, biden, candidate, primary, voters, election, democrats | |
| 18 | protests, insurrection, riots, riot, rioters, peaceful, protestors, protesters, capital, election | |
| 19 | mods, banned, conservatives, post, askconservatives, subreddits, rules, liberal, askaliberal, moderation | |
@@
| 1 | ukraine, russia, war, putin, israel, nato, hamas, iran, are, military | |
| 2 | court, he, case, is, president, supreme, judge, law, trial, immunity | |
| 3 | biden, trump, debate, in, president, vote, has, it, joe, who | |
| 4 | border, immigration, bill, immigrants, asylum, they, mexico, in, is, legal | |
| 5 | kamala, biden, harris, for, they, has, vote, as, debate, if | |
| 6 | left, party, democrats, are, right, republicans, conservatives, liberal, conservative, liberals | |
| 7 | abortion, abortions, birth, women, is, roe, ivf, babies, issue, pregnant | |
| 8 | blue, state, live, california, cities, states, rural, texas, move, liberal | |
| 9 | harris, biden, to, campaign, vote, that, if, they, has, win | |
| 11 | gun, guns, firearms, amendment, shootings, ban, assault, laws, mass, carry | |
| 12 | christian, religion, christians, bible, catholic, christianity, are, islam, commandments, churches | |
| 13 | violence, insurrection, protests, riots, capitol, january, antifa, was, blm, riot | |
| 14 | news, media, cnn, msnbc, the, journalism, propaganda, outlets, bias, journalists | |
| 15 | polls, polling, pollsters, election, betting, trump, is, voters, data, polled | |
| 19 | vote, voting, party, election, volunteer, candidate, blue, in, will, are | |
@@
| 1 | to, is, for, it, as, with, have, woman, just, out | |
| 2 | inflation, economy, prices, that, tax, for, are, trump, tariffs, all | |
| 4 | kamala, to, trump, it, harris, as, biden, have, vote, will | |
| 8 | he, convicted, trump, to, case, trial, judge, jury, prison, court | |
| 9 | men, trans, are, they, it, gender, black, be, woman, as | |
| 13 | border, bill, it, immigrants, in, immigration, wall, are, republicans, asylum |
@@
| 1 | reddit, post, what, re, comments, opinion, google, wrong, downvoted, don | |
| 2 | discs, copies, buy, license, digitally, games, download, store, sales, internet | |
| 3 | platinum, trophies, achievements, plat, platinums, platinumed, getting, hours, enjoy, games | |
| 4 | ssd, storage, 2tb, expansion, hdd, 4tb, heatsink, install, seagate, ps5 | |
| 5 | xbox, gen, console, games, 360, 4k, disc, upgrade, performance, storage | |
| 6 | ps2, ps1, playstation, nes, owned, sega, memories, n64, atari, snes | |
| 7 | xbox, exclusives, microsoft, platform, consoles, sony, market, nintendo, platforms, exclusivity | |
| 8 | gamepass, month, subscription, xbox, games, conversion, buy, gp, price, sales | |
| 9 | he, wife, kids, dad, together, likes, gift, play, happy, roblox | |
| 10 | horizontal, overheating, airflow, cooling, vents, ps5, ventilation, exhaust, cabinet, thermal | |
| 11 | cod, mw3, mw2, warfare, battlefield, warzone, multiplayer, campaigns, modern, bo2 | |
| 12 | sale, price, friday, full, discount, buy, 60, deals, waiting, games | |
| 13 | pc, gaming, ps5, consoles, build, pcs, gpu, performance, windows, steam | |
| 14 | boss, difficulty, fight, level, difficult, enemies, mode, parry, game, hours | |
| 15 | female, she, characters, protagonist, gender, yasuke, white, japanese, samurai, censored | |
| 16 | region, psn, vpn, dlc, accounts, steam, sony, uk, requirement, eu | |
| 17 | router, wifi, ethernet, network, modem, mbps, dns, 5ghz, hotspot, fiber | |
| 18 | drift, stick, controllers, controller, sticks, dualsense, fix, left, launch, joystick | |
| 19 | forza, cars, horizon, turismo, motorsport, motorstorm, arcade, motorfest, racer, sim | |
@@
| 1 | viktor, lore, characters, show, jayce, zaun, ekko, ionia, episodes, hextech | |
| 2 | he, worlds, year, player, top, bad, lck, been, performance, mid | |
| 3 | smurfs, smurfing, accounts, mmr, valve, alt, emerald, ban, matchmaking, level | |
| 4 | dota, friends, life, fun, games, moba, time, addiction, addicted, much | |
| 5 | post, read, comments, argument, opinion, downvotes, downvote, discussion, thread, comprehension | |
| 6 | chat, mute, voice, muting, comms, pinging, toxicity, typing, use, mic | |
| 10 | he, nerfed, nerfs, kit, lane, champ, mid, ult, buff, patch | |
| 11 | support, ult, mid, lane, she, adc, lux, jungle, champion, mage | |
| 12 | fps, gpu, pc, issue, steam, dota, ryzen, reconnect, server, ssd | |
| 15 | baron, draft, they, pick, dragon, game, geng, call, rumble, drakes | |
| 17 | reports, 12k, system, toxic, griefing, communication, overwatch, commends, comm, scores | |
| 18 | mouse, camera, keyboard, hotkeys, buttons, press, cursor, hotkey, screen, bind | |
@@
| 1 | post, comment, conversation, here, argument, don, op, read, have, reply | |
| 2 | abortion, fetus, abortions, woman, choice, medical, unborn, conception, is, care | |
| 4 | wives, roles, equal, church, misogyny, paul, she, feminism, authority, misogynistic | |
| 5 | trans, gender, transgender, dysphoria, identity, intersex, medical, transitioning, cis, transgenderism | |
| 6 | israel, hamas, palestinians, palestine, gaza, genocide, israelis, idf, zionism, conflict | |
| 7 | islam, muslims, muslim, islamic, sharia, religions, islamophobia, europe, quran, iran | |
| 8 | atheist, school, raised, religion, believed, catholic, parents, religious, became, wasn | |
| 9 | slavery, slaves, slave, bible, israelites, exodus, leviticus, servitude, enslaved, testament | |
| 10 | evolution, species, apes, darwin, macroevolution, ancestor, creationism, organisms, scientists, abiogenesis | |
| 11 | science, scientific, scientists, scientist, religion, hypothesis, theories, einstein, physics, newton | |
| 12 | catholics, protestants, catholicism, protestant, pope, orthodoxy, apostolic, protestantism, denominations, bishops | |
| 13 | music, songs, metal, bands, rap, listening, satanic, genre, hymns, taylor | |
| 14 | rapture, end, tribulation, raptured, earth, matthew, generation, time, angels, soon | |
| 15 | hell, torment, place, gehenna, heaven, death, hades, eternity, sheol, soul | |
| 16 | schools, education, teach, teachers, curriculum, oklahoma, commandments, religions, district, classrooms | |
| 18 | dress, clothing, hijab, modesty, woman, dressing, covering, nakedness, muslim, nudity | |
| 19 | religion, religions, people, control, society, world, power, organized, we, masses |
@@
| 1 | israel, hamas, palestinians, palestine, palestinian, israelis, zionist, idf, zionists, zionism | |
| 2 | marry, marriage, relationship, he, married, muslim, family, convert, want, date | |
| 3 | she, muslim, relationship, family, tell, but, will, marry, convert, want | |
| 4 | hijab, women, cover, niqab, clothing, modesty, naked, muslim, she, hijabi | |
| 5 | homosexuality, lgbt, lgbtq, queer, homosexual, homophobic, straight, are, gays, sin | |
| 6 | hindus, indian, bangladesh, caste, hindutva, pakistanis, bangladeshi, bjp, bengali, culture | |
| 7 | woman, rights, islam, feminism, feminist, muslim, misogynistic, misogyny, feminists, is | |
| 9 | sin, forgive, forgiveness, repent, allah, repentance, forgiven, forgives, sinning, repenting | |
| 10 | islam, left, leaving, muslim, started, about, it, believe, convert, didn | |
| 11 | post, troll, account, comments, sub, argument, debate, reply, read, arguments | |
| 12 | universe, god, existence, exist, infinite, evidence, dependent, believe, argument, creation | |
| 13 | fasting, ramadan, eat, water, days, fasts, health, weight, up, if | |
| 14 | alcohol, drink, drinking, addiction, smoking, haram, alcoholic, weed, nicotine, harmful | |
| 15 | arabic, language, translations, quran, read, speak, learning, qur, tafsir, arab | |
| 16 | hadith, hadiths, bukhari, quran, scholars, narration, fabricated, books, authenticity, qur | |
@@
| 1 | phone, media, apps, app, screen, scrolling, reddit, instagram, youtube, tiktok | |
| 2 | porn, addiction, masturbation, is, that, with, sexual, as, have, watching | |
| 3 | weight, eat, diet, and, healthy, protein, foods, calorie, body, meals | |
| 4 | gym, workout, exercise, do, it, week, day, feel, working, like | |
| 5 | adhd, with, medication, diagnosed, can, meds, like, or, get, be | |
| 6 | thank, for, sharing, it, post, to, and, words, hear, advice | |
| 7 | read, books, reading, atomic, life, was, habits, this, changed, helped | |
| 8 | friends, people, social, with, meet, be, talk, group, like, new | |
| 9 | weed, smoking, smoke, quit, nicotine, quitting, smoked, years, turkey, vape | |
| 10 | goals, tasks, task, do, list, it, time, day, work, small | |
| 11 | life, age, young, old, 20s, of, time, are, 40, 30 | |
| 13 | her, relationship, he, with, will, but, ex, love, for, time | |
| 14 | quote, is, man, life, he, be, what, will, quotes, from | |
| 15 | year, was, myself, of, been, 2024, have, years, when, be | |
| 16 | study, studying, exam, hours, time, if, break, exams, this, minutes | |
| 17 | post, comment, this, op, advice, it, re, don, like, understand | |
| 18 | addiction, addictions, drugs, addicted, with, are, can, but, life, will | |
| 19 | women, dating, date, are, be, relationship, will, that, girl, rejection | |
@@
| 1 | her, to, that, the, relationship, is, with, in, do, cheating | |
| 2 | thank, for, much, appreciate, proud, sharing, all, advice, words, keep | |
| 3 | phone, media, app, apps, use, tiktok, and, facebook, scrolling, screen | |
| 6 | ugly, looks, attractive, appearance, and, beauty, women, of, yourself, re | |
| 10 | parents, family, mom, dad, mother, that, he, their, was, be | |
| 19 | thoughts, meditation, mind, thought, thinking, brain, intrusive, mindfulness, think, be |
@@
| 1 | xavi, coach, laporta, manager, season, he, club, team, stay, has | |
| 2 | ref, var, refs, foul, offside, penalty, referee, referees, negreira, madrid | |
| 3 | we, half, game, score, chances, goal, team, goals, defense, up | |
| 4 | ballon, award, vini, won, messi, votes, awards, carvajal, deserved, euros | |
| 5 | mbappe, mbappé, psg, madrid, kylian, ronaldo, player, will, world, and | |
| 6 | vini, vinicius, neymar, him, player, brazil, jr, has, but, fans | |
| 7 | mods, post, comments, sub, reddit, twitter, mod, banned, users, discussion | |
| 9 | kits, logo, shirt, jerseys, nike, authentic, design, crest, replica, badge | |
| 10 | pedri, gavi, fdj, role, midfield, midfielder, injured, pivot, play, gundo | |
| 11 | kroos, modric, toni, luka, modrić, retire, midfield, midfielder, midfielders, season | |
| 12 | contract, sell, clause, salary, million, pay, fee, transfer, loan, value | |
| 16 | yamal, messi, age, 17, kid, talent, him, young, player, will | |
| 18 | madrid, fans, barca, sub, real, support, hate, club, comments, barcelona | |
@@
| 1 | klay, thompson, warriors, was, but, bench, contract, in, with, you | |
| 2 | kerr, minutes, klay, steph, coaching, this, vets, lineups, curry, when | |
| 3 | ad, lebron, center, he, ball, defense, play, can, big, offense | |
| 6 | contract, value, trade, million, salary, pay, cap, player, option, paid | |
| 7 | refs, foul, calls, fouls, officiating, ball, fouled, replay, lakers, bounds | |
| 11 | draft, son, nba, 55th, lebron, bron, his, james, be, lakers | |
| 12 | reddit, post, read, comments, re, opinion, what, said, internet, downvotes | |
| 16 | coach, coaches, coaching, fired, hire, hc, nba, riley, assistants, hiring | |
| 18 | fans, kobe, laker, lakers, lebron, hate, sub, media, haters, hater | |